OpenStudy (anonymous):

solve using subsitution 4x-7y=6 x=10-8y

OpenStudy (anonymous):

4x-7y=6 x=10-8y 4(10-8y)-7y=6 40-32y-7y=6 40-39y=6 -39y=-34 y=34/39 x=10-8(34/39) x=10-(272/39) x=118/39 That should be it